Reports 1-1 of 1 Clear search Modify search
naohisa.satou - 21:30 Tuesday 19 January 2021 (15786) Print this report
Preparation and check of F3
2021.1.19 Sato, Horiuchi

* Moved the F3 into the clean booth.
* Opened the cup of F3.
(The fit of the cup was very tight, so we used a metal spatula, plastic hammer, and alcohol to open it.
* Replaced the FR blade with a reinforced one (JGW-D1706457-2).
* Checked the operation of the FR motor.
Contents of check
1.Motor rotation direction
The direction of rotation of the motor was reversed.
Therefore, the motor wiring was changed as shown below.
With this change, the FR now pulls up the keystone when the right rotation command is given.

[pin No.] [Before] [After]

2. Vertical movement range
Current FR height(set value): 41mm
Upper limit height: 62mm
Lower limit height: 20mm

* Check the resistance of GAS LVDT.
[pin No.] [coil name] [resistance (ohm)]
1-6 Primary 190
2-7 Secondary 75.6
3-8 Actuator 109.6

* Preparation for BF transfer function measurement
In order to prepare for the measurement, the BF keystone was fixed.
The measurement of the transfer function will be done by Miyo-san.

Translated with (free version)
Non-image files attached to this report
Comments to this report:
kouseki.miyo - 12:57 Wednesday 20 January 2021 (15792) Print this report

We measured the noise spectra when we fixed the keystone or not.
Measurement file is /kagra/Dropbox/Measurement/LVDT/GAS/GAS_LVDTs.xml.

According to measurement result, the noise when keystone was fixed is higher than that when keystone was not fixed. Reason of the noise increasing behavior might be height of the keystone because the K1:VIS-ETMY_BF_LVDTINF_GAS_INMON was almost -3000 count which means the keystone was shifted by 2 mm from center position according to plot#1 in klog#15742.

I think we should fix the keystone at several positions to obtain the height dependent noise of GAS_LVDT if we have time.

Images attached to this comment
kouseki.miyo - 15:06 Wednesday 20 January 2021 (15796) Print this report

Mystery was understood. Spectrum when the keystone was fixed is wrong. At this time 2021/01/19 00:00 JST, LVDT calbles are pulled out due to horiday according to Sato-san. So the spectrum shows the ADC like noise...

kouseki.miyo - 10:30 Thursday 21 January 2021 (15803) Print this report

As Sato-san reported in klog#15800, BF was released. So I could compare the LVDT noises in several cases;

  • when the BF keystone was free, which is shown with red line,
  • when the BF keystone was fixed, which is shown with blue dash line,
  • when the cables between the feedthrough and DGS rack were pulled out, which is shown with black dashed line.

Please note that the data taken before January 21 are calibrated with the old value (-0.883 um/count) not new value (-0.848 um/count) as I reported in klog#15788.

Dtt file used for this measurement is saved in "/kagra/Dropbox/Measurements/VIS/LVDT/GAS_LVDTs/2021/0120/GAS_LVDTs_1626.xml".

Images attached to this comment
Search Help